# The Missing Parts: Augmenting Fact Verification with Half Truth Detection

### v1

@@ Appendix B Prompts Used in TRACER
  2. 2.
    Only include assumptions that you are confident in and that serve as a strong basis for the intended conclusion.
(Examples are omitted.)
Claim: {claim}
Intended conclusion: {intention}
Questions:
{questions}
Figure 14: Prompt to generate assumptions.
Prompt:
You are required to do a counterfactual causal inference on a given causal graph.
Argument:
@@
Evaluate $\Delta P(Z\mid\text{do}(\{\text{letter}\}=\neg\{\text{letter}\}))$.
[…]
Options:
A. The probability of $Z$ does not change.
B. The probability of $Z$ increases (Z becomes more likely to be true).
C. The probability of $Z$ decreases (Z becomes less likely to be true).
Please answer with one letter only.
Figure 15: An example of a prompt for evaluating causality, where letter is a placeholder representing the specific assumption to be examined.
A claim may be factually accurate but still misleading due to its implied conclusion. Your task is to refine the veracity assessment of such a claim by considering additional hidden information.
You are given a previously generated fact-checking justification, along with new evidence and an argument supporting the intended conclusion of the claim.
Please determine whether the justification has already addressed the hidden information. Then, refine the veracity of the claim accordingly.
Input:
@@
Instruction: Reassess the veracity of the claim based on the above.
Choose one of the following options (output only the letter):
A. True // B. Half-true // C. False // D. Unverifiable (e.g., the hidden assumption does not support the conclusion, or the information is insufficient)
Your answer (one letter only):
Figure 16: Prompt for reassessment. Here "Evidence" specifies hidden evidence. "Argument" represents the causal graph consisting of intent, claim and assumptions.

### v2

@@ Appendix B Prompt Templates Used in TRACER
2. Only include assumptions that you are confident in and that serve as a strong basis for the intended conclusion.
(Examples are omitted.) Claim: {claim} Intended conclusion: {intention} Questions: {questions}
Prompt: Causality Analysis You are required to do a counterfactual causal inference on a given causal graph. Argument:
@@
Evaluate $\Delta P(Z\mid\text{do}(\{\text{letter}\}=\neg\{\text{letter}\}))$. More specifically, how does the probability of $Z$ change when we set {letter} from {letter} to $\neg${letter}? Options: A. The probability of $Z$ does not change. B. The probability of $Z$ increases (Z becomes more likely to be true). C. The probability of $Z$ decreases (Z becomes less likely to be true). Please answer with one letter only.
Prompt: Re-Assessment A claim may be factually accurate but still misleading due to its implied conclusion. Your task is to refine the veracity assessment of such a claim by considering additional hidden information. You are given a previously generated fact-checking justification, along with new evidence and an argument supporting the intended conclusion of the claim. Please determine whether the justification has already addressed the hidden information. Then, refine the veracity of the claim accordingly. Input:
@@
Instruction: Reassess the veracity of the claim based on the above. Choose one of the following options (output only the letter): A. True // B. Half-true // C. False // D. Unverifiable (e.g., the hidden assumption does not support the conclusion, or the information is insufficient) Your answer (one letter only):
Appendix C Generalization with LLaMA2-7B
Results in Table 11 demonstrate that TRACER yields consistent improvements across metrics when applied to the open-source LLaMA2-7B, with particularly notable gains in Half-True classification.
Model Accuracy Macro-F1 F1(True) F1(Half-True) F1(False)
HiSS 78.3 59.4 44.7 44.4 88.9
HiSS + RA (GPT-3.5-turbo) 81.9 65.7 46.6 60.5 90.1
HiSS + RA (LLaMA2-7B) 82.3 65.4 43.6 61.3 91.2
Table 11: Generalization of TRACER with different backbones.
We further analyze TRACER’s performance across different claim lengths. Using the open-source LLaMA2-7B backbone, we partition test claims into four length ranges by word count. Table 12 shows consistent improvements across all ranges, with larger gains observed for longer claims, which likely offer richer context for intent inference and assumption generation.
Model 4–13 (755) 14–23 (893) 24–34 (308) $\geq$35 (44)
HiSS 80.3 78.5 73.1 75.0
HiSS + RA 83.7 81.8 80.5 81.8
Improvement $\uparrow$3.4 $\uparrow$3.3 $\uparrow$7.5 $\uparrow$6.8
Table 12: TRACER’s performance across different claim lengths (F1 scores) using LLaMA2-7B. Numbers in parentheses indicate the number of examples per length range. Longer claims provide richer context, leading to larger improvements.